Find the surface area generated when the portion of the curve

y2 = 12x from x = 0, x = 3 is rotated about the x-axis.



1
Expert's answer
2021-06-14T15:59:17-0400


If the curve is described as "x=g(y), c\leq y\leq d," then the formula for surface area

becomes


"S=\displaystyle\int_{c}^{d}2\pi y\sqrt{1+(x'_y)^2}dy"


"x=\dfrac{1}{12}y^2=>x'_y=\dfrac{1}{6}y"

"0\leq y\leq 6"

"S=\displaystyle\int_{0}^{6}2\pi y\sqrt{1+(\dfrac{1}{6}y)^2}dy"

"u=1+\dfrac{1}{36}y^2, du=\dfrac{1}{18}ydy"

"\int y\sqrt{1+\dfrac{1}{36}y^2}\ dy=18\int\sqrt{u}du=12u^{3/2}+C"

"=12(1+\dfrac{1}{36}y^2)^{3/2}+C"

"S=24\pi\big[(1+\dfrac{1}{36}y^2)^{3/2}\big]\begin{matrix}
 6 \\
 0
\end{matrix}=24\pi(2\sqrt{2}-1) (units^2)"


"S=24\pi(2\sqrt{2}-1) \text{ square units}"